OpenStudy (sparklestaraa):

Which of the following inequalities is modeled by the graph? x + 4y >= 15; y >= 0 x - 4y >= 15; y >= 0 x + 4y <= 15; y >= 0 -x - 4y >= 15; y >= 0

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (danjs):

I can tell ya that Y is greater than or equal to zero :)

OpenStudy (sparklestaraa):

lol that is In all of them xD @DanJS

OpenStudy (sparklestaraa):

@iLorita

OpenStudy (danjs):

The other line, will be Y is Less than or Equal to something

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (sparklestaraa):

C???

OpenStudy (danjs):

Put each choice into the form \[y \le~~~~~~~or~~~~~y \ge\]first..

OpenStudy (daniellelovee):

plus @DanJS is usually right when helping ppl so you can listen to him/her

OpenStudy (danjs):

\[y \ge \frac{ -1 }{ 4 }x+\frac{ 15 }{ 4 }\] That is the first one, solve the other 3 like that...

OpenStudy (sparklestaraa):

sorry i didnt mean to close ?

OpenStudy (sparklestaraa):

So its B?

OpenStudy (danjs):

Not sure, show me how you decided that and i can check it